# A waveform and time digitization mainboard prototype for the hybrid digital optical module of TRIDENT neutrino experiment

### v2

@@ Abstract
The TRIDENT (Tropical Deep-sea Neutrino Telescope) experiment is a next-generation underwater neutrino observatory planned for deployment in the West Pacific Ocean, designed to detect astrophysical neutrinos through Cherenkov radiation. The full-scale detector will consist of approximately 1000 vertical strings, each equipped with 20 hybrid digital optical modules (hDOMs) containing both photomultiplier tubes (PMTs) and silicon photomultipliers (SiPMs) for comprehensive light detection. This paper presents a custom-designed digitization mainboard prototype for the hDOM, featuring simultaneous 32-channel PMT waveform digitization at 125 MS/s using commercial analog-to-digital converters (ADCs) and 56-channel high-precision time measurement through field-programmable gate array (FPGA)-implemented time-to-digital converters (TDCs). The system demonstrates excellent performance in single photoelectron (SPE) detection with clear pedestal separation, maintains linear response up to 240 PEs, and achieves sub-nanosecond timing resolution for PMT or SiPM pulse edges.
@@ 1 Introduction
Astrophysical neutrinos are unique messengers to study some of the long-standing problems of the universe such as the origins and acceleration mechanism of cosmic rays. The TRopIcal DEep-sea Neutrino Telescope (TRIDENT) [1], planned to be constructed in the West Pacific Ocean, will cover multi-cubic-kilometer of seawaters with around 1000 vertical strings, each around 700 meters long and spaced 70-100 meters apart, adoting a Penrose tiling geometry. Each string will carry 20 hybrid digital optical modules (hDOMs) to detect Cherenkov photons emitted from high-energy charged particles produced in neutrino interactions.
@@
Figure 1 shows the current prototype hDOM. Each hDOM will have a number of photomultiplier tubes (PMTs) for better photon coverage. This also allows local coincidence to reduce backgrounds. Similar approach is taken by the KM3NeT [2], IceCube Upgrade [3], and IceCube-Gen2 experiments [4]. TRIDENT will also use faster-response silicon photomultipliers (SiPMs). The rising time of typical single photoeletron (SPE) signals is about 1 ns [5]. This will improve the photon arrival time measurement of each hDOM and thus the pointing capability of the detector array [5]. However, greater number of channels impose larger challenges for electronics inside the module in terms of limited space (see Figure 1), power consumption budget, cost and so on. KM3NeT chooses to read out the Time-over-threshold (ToT) signals from PMTs and perform digitization of these signals inside an FPGA to measure both the arrival time and length of the ToT signals. For IceCube Upgrade [3], analog-to-digital converters (ADCs) with a sampling rate of 120 MSps are used for waveform digitization. IceCube-Gen2 [4] takes a dual readout approach, signals from each PMT are digitized with a 2-channel ADCs at a sampling rate of 60 MSps.
The typical SPE signal from PMTs used by TRIDENT is characterized by a short negative pulse with 2-3ns rising and 4-5ns falling time (see Figure 2). We would like to have the digitized PMT waveform for possible pulse-shape analyses. A full waveform also gives more precise measurement of photon counts than ToT signals and opens new window for significantly improved efficiency for astrophysical tau neutrino detection [6, 7]. Ideally, to preserve the original pulse shape, ADCs with a sampling rate above several hundreds of MSps are required. However, high sampling rate usually means high power consumptions and cost. Therefore, we aim to have a sampling rate around 100 MS/s. This means we also need a pulse-shaping circuit in front of the ADC to broaden the original PMT pulses. To mitigate the impact on the timing measurements of relatively large sampling interval, the timing information given by the ToT signals from the original pulses can be measured more precisely using time-to-digital converter (TDC) techniques.
@@ 2 The Mainboard Design
Figure 3: The overall design of the mainboard. See text for details.
The design block of the mainboard is shown in Figure 3. It can take at most 32 PMT analog signals. All signal processing circuits for PMT waveform digitization (ADC) and rising edge discrimination, digital signal processing (FPGA), power and clock modules, and so on, are integrated in one board. For this prototype, a 16-channel ADC from ADI (AD9083) is chosen. It supports the JESD204B based high-speed serialized output [8], which helps to accommodate all components and traces in a 22cm-diameter PCB (Figure 4). The ADC is configured so each channel outputs 16-bit resolution digital data (input range of 2 Vpp) with a sampling rate of 125 MSps.
@@
The mainboard integrates a custom-designed White-Rabbit (WR)-based timing module from Sync(Beijing) Technology. This WR module can provide a synchronized clock with sub-ns precision when the mainboard is connected to a upstream WR switch through a small form-factor pluggable plus transceiver (SFP+) port with a fiber optical link. This clock is used as the reference clock for a JESD204B compliant and programmable clock jitter cleaner with internal phased-locked loops (PLL) on the mainboard. This clock cleaner generates necessary clocks for ADC and JESD204B receiver core clock in the FPGA, as well the system reference (SYSREF) signal to achieve deterministic latency among multiple ADCs (the JESD204B subclass 1 system). In addition, several other clocks are required. One clock is needed for configuration (e.g. the PLL and ADCs). Serial transceivers in the FPGA also requires reference clocks (denoted as MGT__\__CLK in Figure 3). One MGT__\__CLK is from the WR module. Other clocks come from an oscillator on the board.
@@
Inside the FPGA, data from ADCs are seperated into channel-by-channel. The way the data are recorded depends on the configured trigger mode. For the self-trigger mode, each channel can be read out independently with baseline suppressed when the incoming data exceed the baseline by a configurable threshold. This mode can be used to record SPE signals. For coincidence trigger mode, a minimum number of channels are required to be simultaneously triggered within a time window. This mode is designed to record physics events while rejecting dark-current signals. For the external trigger mode, all channels of a fixed length are read out upon receiving an external trigger request. This mode is useful for light calibration of PMTs. To record waveform before the trigger, first-in-first-out (FIFOs) in the FPGA are used to cache the data. Recorded ADC data and other information such as the trigger time stamp and channel number are written into separate FIFOs before they are buffered into the DDR3.
@@
In this second circuit, each PMT input signal is fed into a comparator which generates a ToT signal. The threshold of each comparator can be independently adjusted by the output of a DAC. The rising time of each ToT signal is measured by a tapped delay line-based TDC implemented inside the FPGA. Each measurement consists of a coarse counter (with TDC system clock of 4 ns period) and a fine counter which refers to the number of taps delay between the rising edge of the ToT signal and the next nearest rising edge of the system clock. Similarly, the TDC data can be recorded depending on the trigger mode. Together with other information, these data are written into FIFOs and then to the DDR3. A Gigabit Ethernet protocol SiTCP [9] is used to transfer these data from the FPGA to the DAQ server through another SFP+. In the future it might be possible to use the same WR link to transmit data to reduce the number of optical fibers along each string. This is still under development.
@@
Besides the above-mentioned functionalities, the mainboard is also designed to take at most 24 LVDS ToT signals from the Samtec high-speed connector shown in Fig. 4. These signals come from another daughter card where each SiPM analog output is discriminated against a threshold configured by the mainboard. Therefore, we can measure the ToT signals from SiPMs using the same TDC techniques in the FPGA. Another 1 MSps ADC (slow ADC in Figure 3) is used for digitizing the acoustic signal. An accelerometer gyroscope is used to monitor the angle, acceleration, angular velocity, magnetic field and temperature. Finally, this board has a 14-pin connector, which can be used to control and monitor the HV of each PMT through a separate HV control board.
@@ 3 Performance
In the following, we report the performance of the charge measurement from the digitized waveform using ADCs, including the SPE signal and the linear range, which are among the most important criteria for waveform digitization. Afterwards, we report the resolution of the SPE rising-edge time measurement using the TDC implemented in the FPGA.
@@ 3.1 Charge Measurement
A laser source is used to evaluate the performance of the waveform digitization, shown in Figure 5. Pulsed laser lights (typical width of 10 ps) are injected into the PMT together with a trigger request into the mainboard. The light intensity is controlled with optical filters. At each intensity, we measure the charge from a larger number of events. We first tune the intensity for SPE calibration. The left panel of Figure 6 shows the measured charge distribution by integrating the digitized waveform in a fixed time interval around the pulse. The pedestal peak exhibits an RMS of 0.06 PE, coming from the integrated electronic noise. In contrast, the SPE signal peak shows an RMS of 0.4 PE, dominated by the statistical fluctuations in the PMT gain. This indicates that the noise contribution becomes negligible for charge measurements of the PMT’s smallest signals, as the SPE resolution is overwhelmingly determined by gain variation. The right panel of Figure 6 shows a typical waveform of the SPE signals recorded by the mainboard. The intensity of the laser source is then tuned. Figure 7 shows the measured charged distribution and a typical waveform at the largest intensity of the laser source.
@@
Figure 7: Left, the measured charge distribution with a high-intensity light source. Right, a typical waveform recorded at this light-intensity.
Figure 8: The linearity test result of the mainboard with a PMT and a light source. Reference measurement is from an oscilloscope. The dashed line indicates ideal linearity.
To evaluate the linearity performance of the mainboard, we compare the charge measured from the mainboard with that from an oscilloscope with a sampling rate of 10 GS/s, because the intensity of laser source is not calibrated. Figure 8 shows that our mainboard has good linearity up to 240 PE. For comparison, IceCube collaboration [3] reported a linear dynamic range of 150 PE with their DOM mainboard for the Upgrade experiment. IceCube-Gen2 collaboration [4] uses dual-readout design for each PMT (high gain with anode and low gain with eighth dynode), achieving a linear dynamic range of 40 PE (high gain) and 2500 PE (low gain).
@@ 3.2 Timing Measurement
As mentioned above, we also want to measure the rising edge time of the ToT signal of each PMT channel with the delay chains inside the FPGA. For each TDC channel in the FPGA, the delay of each element along the delay line is not identical, but can be measured from the distribution of the fine counter by collecting a large amount of events which are asynchronized with the TDC system clock. This is the so-called bin-by-bin calibration or statistical code density test. We can use SPE events to perform this calibration in-situ.
In addition, signal propagation delay from the input port to the first element in the FPGA varies among different channels. These channel-by-channel variations can be calibrated using events with the same arrival time. In the lab, we can use a signal-splitter board to distribute the same pulse with equal-length cables to 32 input ports. In situ, this can be calibrated using the Potassium-40 backgrounds from the seawater. This in-situ calibration can also remove possible systematic transit-time variations among different PMTs.
Figure 9: The distribution of the time difference measured by two TDC channels after calibrations using identical SPE-like pulses, from a pulse generator that produces double-PE-like pulses which are then splitted equally.
To evaluate the timing resolution of the TDC, we inject two identical SPE-like pulses into two channels. The measured time difference is shown in Figure 9. The RMS is much smaller than the transit time spread (TTS) of the PMTs which is at the order of ns. Therefore, in situ we expect to achieve a timing precision of O(ns) for PMTs. This mainboard can also measure at most 24 ToT signals from SiPM arrays using the same FPGA-based TDC technique. Therefore, with SiPMs, even better timing precision can be achieved. Previous study [5] shows that single photon time resolution with an array of $4\times 44 × 4$ SiPMs is as low as 300 ps FWHM, which is much smaller than the TTS of PMT. Notably, the IceCube-gen2 collaboration [4] demonstrated that PMT digitized pulses can be fitted with a smooth curve, achieving nanosecond-level time resolution. In contrast, our TDC implementation enables direct hardware-level time measurement.

### v3

@@ Abstract
The TRIDENT (Tropical Deep-sea Neutrino Telescope) experiment is a next-generation underwater neutrino observatory planned for deployment in the West Pacific Ocean, designed to detect astrophysical neutrinos through Cherenkov radiation. The full-scale detector will consist of approximately 1000 vertical strings, each equipped with 20 hybrid digital optical modules (hDOMs) containing both photomultiplier tubes (PMTs) and silicon photomultipliers (SiPMs) for comprehensive light detection. This paper presents a custom-designed digitization mainboard prototype for the hDOM, featuring simultaneous 32-channel PMT waveform digitization at 125 MS/s using commercial analog-to-digital converters and 56-channel high-precision time measurement through field-programmable gate array -implemented time-to-digital converters. The system demonstrates excellent performance in single photoelectron (PE) detection with clear pedestal separation, maintains linear response up to 240 PEs, and achieves sub-nanosecond timing resolution for PMT or SiPM pulse edges.
@@ 1 Introduction
Astrophysical neutrinos are unique messengers to study some of the long-standing problems of the universe such as the origins and acceleration mechanism of cosmic rays. The TRopIcal DEep-sea Neutrino Telescope (TRIDENT) [TRIDENT:2022hql], planned to be constructed in the West Pacific Ocean, will cover multi-cubic-kilometer of seawaters with around 1000 vertical strings, each around 700 meters long and spaced 70-100 meters apart, adopting a Penrose tiling geometry. Each string will carry 20 hybrid digital optical modules (hDOMs) to detect Cherenkov photons emitted from high-energy charged particles produced in neutrino interactions.
@@
Figure 1 shows the current prototype hDOM. Each hDOM will have a number of photomultiplier tubes (PMTs) for better photon coverage. This also allows local coincidence to reduce backgrounds. Similar approach is taken by the KM3NeT [KM3NeT:2019shc], IceCube Upgrade [IceCube:2023upt], and IceCube-Gen2 experiments [IceCube-Gen2:2023rji]. TRIDENT will also use faster-response silicon photomultipliers (SiPMs). The rising time of typical single photoelectron (SPE) signals is about 1 ns [Zhi:2024dmr]. This will improve the photon arrival time measurement of each hDOM and thus the pointing capability of the detector array [Zhi:2024dmr]. However, greater number of channels impose larger challenges for electronics inside the module in terms of limited space (see Figure 1), power consumption budget, cost and so on. KM3NeT chooses to read out the time-over-threshold (ToT) signals from PMTs and perform digitization of these signals inside an FPGA to measure both the arrival time and length of the ToT signals. For IceCube Upgrade [IceCube:2023upt], analog-to-digital converters (ADCs) with a sampling rate of 120 MS/s are used for waveform digitization. IceCube-Gen2 [IceCube-Gen2:2023rji] takes a dual readout approach; signals from each PMT are digitized with a 2-channel ADCs at a sampling rate of 60 MS/s.
The typical SPE signal from PMTs used by TRIDENT is characterized by a short negative pulse with 2-3ns rising and 4-5ns falling time (see Figure 2). We would like to have the digitized PMT waveform for possible pulse-shape analyses. A full waveform also gives more precise measurement of photon counts than ToT signals and opens new window for significantly improved efficiency for astrophysical tau neutrino detection [Tian:2023+9, Tian:PhD]. Ideally, to preserve the original pulse shape, ADCs with a sampling rate above several hundreds of MS/s are required. However, high sampling rate usually means high power consumptions and cost. Therefore, we aim to have a sampling rate around 100 MS/s. This means we also need a pulse-shaping circuit in front of the ADC to broaden the original PMT pulses. To mitigate the impact on the timing measurements of relatively large sampling interval, the timing information given by the ToT signals from the original pulses can be measured more precisely using time-to-digital converter (TDC) techniques.
@@ 2 The Mainboard Design
Figure 3: The overall design of the mainboard. Some of the main parts are: ADC (ADI AD9083), FPGA (Xilinx XC7k325T), DDR3 (Mircon MT41J128M16JT), Comparator (TI LTV3603), and PLL (TI LMK04610).
The design block of the mainboard is shown in Figure 3. It can take at most 32 PMT analog signals. All signal processing circuits for PMT waveform digitization (ADC) and rising edge discrimination (TI LTV3603), digital signal processing (FPGA, Xilinx XC7k325T), power and clock modules (TI LMK04610), and so on, are integrated in one board. For this prototype, a 16-channel ADC from ADI (AD9083) is chosen. It supports the JESD204B based high-speed serialized output [jedec], which helps to accommodate all components and traces in a 22-cm-diameter PCB (Figure 4). The ADC is configured, so each channel outputs 16-bit resolution digital data (input range of 2 Vpp) with a sampling rate of 125 MS/s.
@@
Parameter or fuctionality Description
Number of ADC channels 32 instantaneous sampling at 125MS/s for PMT signals, and 1 channel at 1 MS/s for acoustic signal
Number of TDC channels 32 for PMT and 24 for SiPM
Clock synchronization WR integrated
Trigger mode Self-trigger, Coincidence trigger or External trigger
PMT High voltage Control and monitor
Accelerometer gyroscope Monitor the angle, acceleration and so on
Table 1: Key design parameters or functionalities of the mainboard.
Table 1 shows the key design parameters or functionalities of the mainboard. In this paper, we focus on the waveform and time digitization of the PMT signals.
Each input PMT signal is processed by two analog front-end circuits in parallel. One is the signal conditioning circuit for ADCs (Figure 5), another one is the comparator circuit for TDCs. In the first circuit, the single-ended input signal is first processed with two preamplifiers and low pass RC filters. The first preamplifier is mainly for impedance matching of 50 $\Omega$ and configured with unit gain. But due to its limited bandwidth, it can introduce small broadening effect. The two RC circuits perform the main broadening of the pulse shape. The second amplifier is set to be unit gain (redundant design for this prototype). Then it is converted into a differential pair through a differential amplifier. The gain of this amplifier is set to be two. Meanwhile, the baseline level is adjusted to $\sim$0.9V to properly use the $\pm 1$V dynamic range of the ADC. To illustrate the effect of this circuit, for SPE signal, the rising edge and falling edge are enlarged to about 15 ns and 30 ns, respectively. The amplitude is reduced to approximately 7 mV. The digital data from ADC are transferred to the FPGA for online processing.
@@
The mainboard integrates a custom-designed White-Rabbit (WR)-based timing module from Sync(Beijing) Technology. This WR module can provide a synchronized clock with sub-ns precision when the mainboard is connected to a upstream WR switch through a small form-factor pluggable plus transceiver (SFP+) port with a fiber optical link. This clock is used as the reference clock for a JESD204B compliant and programmable clock jitter cleaner with internal phased-locked loops (PLL) on the mainboard. This clock cleaner generates necessary clocks for ADC and JESD204B receiver core clock in the FPGA, as well the system reference (SYSREF) signal to achieve deterministic latency among multiple ADCs (the JESD204B subclass 1 system). In addition, several other clocks are required. One clock is needed for configuration (e.g. the PLL and ADCs). Serial transceivers in the FPGA also require reference clocks (denoted as MGT_\_CLK in Figure 3). One MGT_\_CLK is from the WR module. Other clocks come from an oscillator on the board.
@@
Inside the FPGA, data from ADCs are separated into channel-by-channel. The way the data are recorded depends on the configured trigger mode. For the self-trigger mode, each channel can be read out independently with baseline suppressed when the incoming data exceed the baseline by a configurable threshold. This mode can be used to record SPE signals. For coincidence trigger mode, a minimum number of channels are required to be simultaneously triggered within a time window. This mode is designed to record physics events while rejecting dark-current signals. For the external trigger mode, all channels of a fixed length are read out upon receiving an external trigger request. This mode is useful for light calibration of PMTs. To record waveform before the trigger, first-in-first-out (FIFOs) in the FPGA are used to cache the data. Recorded ADC data and other information such as the trigger time stamp and channel number are written into separate FIFOs before they are buffered into the DDR3 (Mircon MT41J128M16JT).
@@
In this second circuit, each PMT input signal is fed into a comparator which generates a ToT signal. The threshold of each comparator can be independently adjusted by the output of a DAC. The rising time of each ToT signal is measured by a tapped delay line-based TDC implemented inside the FPGA, using the CARRY4 block, each of which has 4 delay taps. In this work, each TDC line consists of 96 CARRY4s, spanning two clock regions. Each measurement consists of a coarse counter (with TDC system clock of 4 ns period) and a fine counter which refers to the number of taps delay between the rising edge of the ToT signal and the next nearest rising edge of the system clock. One unit of fine counter corresponds to 12 ps on average, which limits the intrinsic precision of the TDC in FPGA. Similarly, the TDC data can be recorded depending on the trigger mode. Together with other information, these data are written into FIFOs and then to the DDR3. A Gigabit Ethernet protocol SiTCP [4545224] is used to transfer these data from the FPGA to the DAQ server through another SFP+. In future, it might be possible to use the same WR link to transmit data to reduce the number of optical fibers along each string. This is still under development.
@@
Besides the above-mentioned functionalities, the mainboard is also designed to take at most 24 LVDS ToT signals from the Samtec high-speed connector shown in Fig. 4. These signals come from another daughter card where each SiPM analog output is discriminated against a threshold configured by the mainboard. Therefore, we can measure the ToT signals from SiPMs using the same TDC techniques in the FPGA. Another 1 MS/s ADC (slow ADC in Figure 3) is used for digitizing the acoustic signal. An accelerometer gyroscope is used to monitor the angle, acceleration, angular velocity, magnetic field and temperature. Finally, this board has a 14-pin connector, which can be used to control and monitor the HV of each PMT through a separate HV control board.
@@ 3 Performance
In this design, ADC plays an important role. But it should be emphasized any ADC performance might be affected by the front-end amplifier circuits. Particularly, in our case, we have three amplifiers. For dynamic performance, one major specification is the effective number of bits (ENOB). To illustrate the performance, we use a 1.13 MHz sin-wave ($\sim 0.9$ full scale) as input. An FFT analysis (Figure 6) shows the ENOB is 9.7. Major static performance parameters include differential nonlinearity (DNL) and integrated nonlinearity (INL). However, for the chosen AD9083 ADC, these parameters are hard to interpret due to its first order sigma-delta architecture [forum].
Figure 6: Frequency domain representation of the digitized samples with 1.13 MHz sin-wave input.
In this paper, we focus on the experimental performance of the mainboard related to PMT signals. We report the performance of the charge measurement from the digitized waveform using ADCs, including the SPE signal and the linear range, which are among the most important system-level criteria for waveform digitization. Afterward, we report the resolution of the SPE rising-edge time measurement using the TDC implemented in the FPGA.
@@ 3.1 Charge Measurement
A laser source is used to evaluate the performance of the waveform digitization, shown in Figure 7. Pulsed laser lights (typical width of 10 ps) are injected into the PMT together with a trigger request into the mainboard. The light intensity is controlled with optical filters. At each intensity, we measure the charge from a larger number of events. We first tune the intensity for SPE calibration. The left panel of Figure 8 shows the measured charge distribution by integrating the digitized waveform in a fixed time interval around the pulse. The pedestal peak exhibits an RMS of 0.06 PE, coming from the integrated electronic noise. In contrast, the SPE signal peak shows an RMS of 0.4 PE, dominated by the statistical fluctuations in the PMT gain. This indicates that the noise contribution becomes negligible for charge measurements of the PMT’s smallest signals, as the SPE resolution is overwhelmingly determined by gain variation. At the same intensity, the PMT pulses are recorded by an oscilloscope with a sampling rate of 10 GS/s. The noise performance is much better in the oscilloscope, however, the SPE peak is almost the same compared to the mainboard.
The right panel of Figure 8 shows a typical waveform of the SPE signals recorded by the mainboard. The intensity of the laser source is then tuned. Figure 9 shows the measured charged distribution and a typical waveform at the largest intensity of the laser source.
@@
Figure 9: Left, the measured charge distribution with a high-intensity light source. Right, a typical waveform recorded at this light intensity.
Figure 10: The linearity test result of the mainboard with a PMT and a light source. Reference measurement is from an oscilloscope. The data are fitted with a linear curve $y=p0+p1\times x$.
To evaluate the linearity performance of the mainboard, we compare the charge measured from the mainboard with that from the oscilloscope, because the intensity of laser source is not calibrated. Figure 10 shows that our mainboard has good linearity up to 240 PE. The data are fitted with a straight line, the obtained R-squared value is 0.999. At all light intensities, the relative deviations of the mainboard’s measurements compared to the oscilloscope are within 10%. Overall, the mainboard’s results are slightly lower. For comparison, IceCube collaboration [IceCube:2023upt] reported a linear dynamic range of 150 PE with their DOM mainboard for the Upgrade experiment. IceCube-Gen2 collaboration [IceCube-Gen2:2023rji] uses dual-readout design for each PMT (high gain with anode and low gain with eighth dynode), achieving a linear dynamic range of 40 PE (high gain) and 2500 PE (low gain).
@@ 3.2 Timing measurement
As mentioned above, we also want to measure the rising edge time of the ToT signal of each PMT channel with the delay chains inside the FPGA. For each TDC channel in the FPGA, the delay of each element along the delay line is not identical, but can be measured from the distribution of the fine counter by collecting a large amount of events which are asynchronized with the TDC system clock. This is the so-called bin-by-bin calibration or statistical code density test. We can use SPE events to perform this calibration in-situ. The average delay is measured to be 12 ps (denoted as LSB of the TDC). As shown in Figure 11, the relative difference of each delay to the average, namely the DNL, ranges from -1 LSB to 2 LSB, and the INL ranges from -3 LSB to 10 LSB. To estimate the intrinsic time resolution of the TDC, two identical digital pulses are directly sent to two TDC delay lines via two SMA connectors on board. The RMS of the time difference is about 12 ps.
Figure 11: The DNL (left) and INL (right) of the TDC. The LSB refers to the average delay of 12 ps. The feature around fine counter of 180 is caused by the two clock region boundaries in the TDC delay line.
In addition, signal propagation delay from the input port to the first element in the FPGA varies among different channels. These channel-by-channel variations can be calibrated using events with the same arrival time. In the laboratory, we can use a signal-splitter board to distribute the same pulse with equal-length cables to 32 input ports. In situ, this can be calibrated using the Potassium-40 backgrounds from the seawater. This in situ calibration can also remove possible systematic transit-time variations among different PMTs.
Figure 12: The distribution of the time difference measured by two TDC channels after calibrations using identical SPE-like pulses, from a pulse generator that produces double-PE-like pulses which are then split equally.
To evaluate the timing resolution of the TDC for comparator’s output, we inject two identical SPE-like pulses into two channels. The measured time difference is shown in Figure 12. The RMS is much smaller than the transit time spread (TTS) of the PMTs which is at the order of ns. Therefore, in situ we expect to achieve a timing precision of O(ns) for PMTs. This mainboard can also measure at most 24 ToT signals from SiPM arrays using the same FPGA-based TDC technique. Therefore, with SiPMs, even better timing precision can be achieved. Previous study [Zhi:2024dmr] shows that single photon time resolution with an array of $4\times 4$ SiPMs is as low as 300 ps FWHM, which is much smaller than the TTS of PMT. Notably, the IceCube-gen2 collaboration [IceCube-Gen2:2023rji] demonstrated that PMT digitized pulses can be fitted with a smooth curve, achieving nanosecond-level time resolution. In contrast, our TDC implementation enables direct hardware-level time measurement.